# CI/CD Configuration Practices in Open-Source Android Apps: An Empirical Study

### v1

@@ 9 FDroid 4 0 0 1 5
Deployment practices vary among developers. Notably, 118 (50%) projects use GitHub tools, such as Releases and Pages, while 52 (22%) projects use external scripts (e.g., shell scripts). We also observe that Fastlane is used by 44 (18%) projects, whereas only a few projects are found to be deployed on unofficial platforms, such as F-Droid (5 projects). These results suggest that developers may face challenges in automating app deployments to the Google Play Store, as the most popular GitHub Actions template recommends manually uploading an APK or AAB file via the Play Console [^14] https://github.com/marketplace/actions/upload-android-release-to-play-store. However, deployment through the console requires APKs to be signed by the developer’s private key, which ensures developer’s authenticity, but still adds an extra layer of complexity to the process. This highlights the need for better automation solutions to deploy to the Google Play Store directly through GitHub Actions. Travis CI projects rarely automate direct deployments to the Google Play Store–only one project does. Instead, 91 out of 105 projects use other services, such as GitHub Pages, likely to host documentation, release notes, or web-based demos of their apps, revealing distinct deployment habits compared to projects that adopt GitHub Actions. For CircleCI, most users face hurdles and resort to third-party solutions like Fastlane, with only 18 projects including deployment automation and three deploying directly to the Google Play Store. GitLab CI/CD users automate deployments in 13 projects, mainly using Fastlane (five projects), as well as AWS S3, external scripts, and SSH for web hosting. Only two projects deploy directly to the Google Play Store.
@@
Listing 11 provides two examples of how deployment tools are adopted in CI/CD. One example uses Fastlane with Travis CI, while the other showcases direct deployment to the Google Play Store with GitHub Actions. These examples illustrate the complexity of configuring deployments in CI/CD environments, given the extensive number of parameters involved. Moreover, some of these tools are proprietary and may require payment. This indicates diverse deployment preferences and challenges across CI/CD services, suggesting the need for more integrated and automated solutions to streamline mobile app deployments to popular platforms like the Google Play Store.
@@
Summary of RQ 1 CI/CD configurations are not standardized and vary in complexity across categories and CI/CD services. Most efforts focus on pipeline setup, comprising over 81% of directives. Testing is performed in nearly half of the projects, with a predominant focus on standard unit testing, whereas deployment is infrequent, occurring in only 9% of the projects.
@@ 4.2. RQ2: How do CI/CD configurations evolve in Android apps?
In this RQ, we collect and process the commits in which developers change CI/CD configuration files to understand the change complexity, the evolution trends, and the factors associated with change frequency.
Analyzing complexity and evolution of changes. To analyze the evolution of CI/CD configuration complexity, we segment the data by project and commit time period (monthly). To visualize the number of changes (commits) per project across CI/CD services, we use boxplots in which we display commit distribution overall and per CI/CD services, showing variations and outliers. To analyze the evolution of changes across various CI/CD services over the years, we start by identifying the first commit for each project to grasp the initial configuration state. We then calculate the total number of changing commits for each project. These counts are normalized by the total number of active projects adopting CI/CD within a given year, accounting for the newly adopting projects, as follows:
@@
Understanding factors associated with CI/CD change frequency. To understand the factors associated with CI/CD maintenance activity in software projects, by identifying the significant variables associated with the number of configuration commits in the studied projects using linear regression. To do this, we fit an Ordinary Least Squares (OLS) model, a common method for estimating the parameters in a linear regression model identifying the relationships between variables. To fit our model, we use common project characteristics provided by GitHub about software repositories shown in Table 6. Specifically, we use the number of configuration commits as the dependent variable and the other project characteristics as independent variables. We assess the performance of the model using the adjusted $R^{2}$, which measures the proportion of variance in the dependent variable explained by the independent variables. The adjusted $R^{2}$ helps to adjust the obtained $R^{2}$ value for the number of predictors, thus preventing overfitting when using multiple variables. An $R^{2}$ value closer to $1.00$ indicates a better fit. We also evaluate the significance of each factor based on the p-values produced by the model, and identify the directionality of their relationships through coefficient analysis (i.e., either positively associated ($\nearrow$) or negatively associated ($\searrow$)).
@@
Observation 2.2: The majority of projects (52%) show a decreasing trend in CI/CD changes (commits) over time, with 28% and 20% exhibit an increasing or steady rate, respectively, of actual configuration lines modified, suggesting simplification or stabilization in CI/CD adoption. Figure 5 shows an overall trend of configuration changes (commits) over time across the different CI/CD services. We observe that GitHub Actions shows a rapid increase in configuration changes, due to shifting towards it after its introduction in 2018. In contrast, Travis CI, exhibits a decline after 2017, nearly reaching zero configuration changes after 2020, likely due to the decreased adoption among GitHub projects due to being less popular for open-source projects. CircleCI and GitLab CI/CD demonstrate steady growth peaking in 2018-2019, followed by a steady decrease, indicating moderate user retention, likely amidst the introduction of GitHub Actions. However, the surge in configuration commits for projects adopting GitHub Actions may reflect more than mere popularity; it could also suggest potential instability or complexity, prompting frequent adjustments and updates by developers.
@@
Figure 5. Boxplots showing the average number of configuration changes (commits) over time across CI/CD services
With respect to the complexity of configuration changes, we find conflicting trends in the addition and deletion of directives. Specifically, we observe 19% of projects with an increasing number of changes, 42% with a decreasing number, and 41% with no significant change. This indicates that many projects are either simplifying their CI/CD setups or maintaining consistent configuration complexity. These trends highlight a dynamic CI/CD landscape, where some projects become more complex or streamlined while others remain consistent, reflecting differing strategies and stages in project lifecycles.
Moreover, we observe that the frequency of commits has a very weak positive correlation with the number of lines added and deleted, with Pearson coefficients of 0.035 and 0.075, respectively. This indicates no strong link between commit frequency and commit complexity. However, there is a moderate correlation between the trends of committed changes and lines added/deleted over time (Pearson coefficients of 0.27 and 0.28, respectively). This suggests a general increase in development activity and complexity, with commit frequency moderately linked to code changes.
Observation 2.3: On average, CI/CD configurations undergo bi-monthly maintenance (61 days), yet many projects have more frequent configuration changes (a median of 2.8 hours), revealing diverse CI/CD maintenance practices. Our analysis of intervals between CI/CD configuration changes across various projects reveals a notable disparity: the average time between changes is approximately 61 days, suggesting a bi-monthly update frequency, while the median of about 2.8 hours indicates that many projects have much more frequent updates. This significant difference, coupled with a high standard deviation of about 84 days, suggests a wide range of maintenance practices in CI/CD configurations. This variation reflects the diverse approaches to CI/CD management among different projects, influenced by factors such as project requirements, team dynamics, and development methodologies.
@@
Observation 2.5: Active issue tracking, repository size/age, and community engagement are significantly associated with CI/CD configuration maintenance. Table 8 presents the results of the OLS linear regression model, expressed by the modeling coefficients, p-values, significance, and relationship direction. Though our model obtains an adjusted $R^{2}$ value of $0.139$, meaning it explains about 14% of variability in the number of configuration commits, it still offers valuable insights into some of the key factors associated with CI/CD pipeline maintenance activity. We observe that the number of reported issues is significantly associated with CI/CD commit frequency, indicating that developers are likely to update CI/CD pipelines to address issues in their projects. Similarly, repository size and age positively correlate with configuration commits, suggesting larger and older repositories attract more CI/CD activity, possibly due to a larger user base or more bugs/features. Furthermore, we find that repository stars and watchers also positively correlate with commit frequency related to CI/CD configurations, making popular projects more active in addressing issues and deploying app updates, which highlights community engagement’s role in development activity. Unlike stars and watchers, we find that repository forks are inversely associated with commit frequency. Forks often represent independent development paths, leading to fewer contributions to the original CI/CD pipelines, thus reducing centralized commit activity. Our analysis also explores the relationship between CI/CD services and maintenance activity. We find that integrating GitHub Actions is a significant positive contributor to commit frequency. Other CI/CD services might have positive relationships with commit activity, but lack statistical significance, indicating that evidence is not robust enough to reliably confirm their association. Overall, this analysis highlights the importance of certain factors, such as issue tracking, repository age, community engagement, and adoption of CI/CD practices, in fostering an active development environment, while also identifying potential deterrents to commit activity.
@@
Summary of RQ 2 Developers regularly change CI/CD pipelines, primarily focusing on setup and build phases, with most projects showing a trend toward simplification over time. While maintenance likely occurs bi-monthly on average, many projects update CI/CD pipelines more often, especially those using GitHub Actions. Increased maintenance activity correlates significantly with active issue tracking, project size/age, and community engagement.
@@ 4.3. RQ3: What are the common themes in CI/CD maintenance activities in Android apps?
Manual thematic analysis of commit messages. We employ open coding to perform a thorough manual analysis of a statistically significant and stratified sample of $384$ commit messages. This sample size is based on a $95\%$ confidence level and a ±$\pm 5$ confidence interval. Our goal in this manual analysis is to identify recurring themes and categories that explain the reasons for CI/CD configuration changes. Two co-authors, with expertise in software engineering and CI/CD, participate in the manual thematic analysis of the commit messages. They engage in collaborative discussion sessions to ensure a unified interpretation and categorization. When any ambiguity or disagreement arises in the coding of messages or issues, the third co-author is consulted to achieve consensus and make a final decision, thereby ensuring the robustness and reliability of our categorization process.
Automated analysis of commit messages. To further validate the results of our manual analysis, we adopt Latent Dirichlet Allocation (LDA) as a supplementary step to verify the themes identified manually. Recognizing the importance of parameter tuning in LDA for obtaining meaningful and reliable topic distributions, we follow the recommendations provided by Agrawal et al. (2018). Specifically, we employ LDA with Differential Evolution (LDADE) for parameter optimization, focusing on determining the optimal number of topics and tuning the alpha and beta parameters accordingly. As a result, we obtain alpha and beta values of $0.97$ and $0.73$, respectively, and identify $10$ as the optimal number of topics. Moreover, before performing LDA, we preprocess commit messages to ensure their suitability for topic modeling. First, URLs, email addresses, and default GitHub commit messages are removed. Then, stop words, numbers, and common abbreviations are filtered out, focusing on meaningful content. Lemmatization is also performed for each commit to reduce words to their root forms, and non-English characters are excluded. Also, very short messages (i.e., having fewer than three characters) are discarded as they are often non-informative and do not provide enough context. This thorough preprocessing ensures the text data is clean, standardized, and ready for effective topic modeling with LDA. Finally, we utilize the optimized parameters and processed data to perform LDA topic modeling. For each identified topic, we extract the eight main associated keywords, assign each commit message to the topic with the highest probability of occurrence, calculate the percentage of commit messages that belong to each topic, and assess their relevance to the manually identified themes.
@@
Table 9 lists the common themes of the commit messages identified by our open coding analysis. We describe our identified themes in the following. Our analysis of open codes derived from our manual analysis of commit messages in CI/CD configurations of Android applications reveals several key insights into the current practices and trends in this area.
@@
Observation 3.2: Over a third of configuration changes are focused on updating the build environment, tools, and dependencies. Our results shows that CI build configuration are more frequently change for the sake for Environment Update (14.7%), Tool Update (12.6%), and Dependency Update (9.4%). Environment updates are particularly crucial as they demonstrate the necessity of maintaining current environments to align with technological advancements and security protocols. By consistently incorporating these updates, mobile apps remain compatible with the latest technologies, mitigating vulnerabilities inherent in older systems. Moreover, tool and dependency updates highlight the essential need for ongoing management and upgrades of CI/CD pipelines. Regular updates ensure efficient development processes, while managed dependencies keep mobile app components functional and compatible, leading to fewer interruptions and more reliable software performance. While developers should stay updated on CI/CD trends to keep their mobile apps robust, CI/CD service providers should also ensure dependencies are secure (Gruhn et al., 2013) and suggest necessary updates.
@@
Observation 3.3: Testing Improvement and Deployment Improvement messages account for 12.8% and 8.4%, respectively. Robust testing and regression in CI/CD pipelines ensures early detection and resolution of bugs, leading to the development of more reliable and high-quality mobile apps. Similarly, deployment ensures rapid and reliable and efficient release cycles, thus reducing the time-to-market for mobile app updates. Despite the importance of frequently maintaining testing and deployment, our results show relatively fewer changes to them compared to other practices, which could be attributed to several factors. These processes might already be well-established and mature within the organization, hence requiring fewer changes and improvements over time, such as enable more features (e.g., add test coverage and add deployment to playstore) or fixing minor relevant issues (e.g. try number two for multiple devices and fix artifact name). In some cases, testing and deployment might not be considered immediate strategic priorities for developers compared to other areas such as feature development, user experience, or performance optimization (e.g., committing to fix tests later). Still, developers are advised to continuously monitor testing and deployment processes, given the critical importance of robust testing and deployment in mobile app deployment, making ongoing improvements and updates are essential to adapt to new technologies and changing requirements. Likewise, researchers should pay more attention to developing techniques that recommend potential co-changes in testing or deployment configuration in response to changes in source code.

### v2

@@ 10 Semantic-release 2 0 0 2 4
Deployment practices vary among developers. Notably, 118 (50%) projects use GitHub tools, such as Releases and Pages, while 52 (22%) projects use external scripts (e.g., shell scripts). We also observe that Fastlane is used by 44 (18%) projects, whereas only a few projects are found to be deployed on unofficial platforms, such as F-Droid (5 projects). These results suggest that developers may face challenges in automating app deployments to the Google Play Store, as the most popular GitHub Actions template recommends manually uploading an APK or AAB file via the Play Console [^25] https://github.com/marketplace/actions/upload-android-release-to-play-store. However, deployment through the console requires APKs to be signed by the developer’s private key, which ensures developer’s authenticity, but still adds an extra layer of complexity to the process. This highlights the need for better automation solutions to deploy to the Google Play Store directly through GitHub Actions. Travis CI projects rarely automate direct deployments to the Google Play Store—only one project does. Instead, 91 out of 105 projects use other services, such as GitHub Pages, likely to host documentation, release notes, or web-based demos of their apps, revealing distinct deployment habits compared to projects that adopt GitHub Actions. For CircleCI, most users face hurdles and resort to third-party solutions like Fastlane, with only 18 projects including deployment automation and three deploying directly to the Google Play Store. GitLab CI/CD users automate deployments in 13 projects, mainly using Fastlane (five projects), as well as AWS S3, external scripts, and SSH for web hosting. Only two projects deploy directly to the Google Play Store.
@@
Listing 11 provides two examples of how deployment tools are adopted in CI/CD. One example uses Fastlane with Travis CI, while the other showcases direct deployment to the Google Play Store with GitHub Actions. These examples illustrate the complexity of configuring deployments in CI/CD environments, given the large number of parameters involved. Moreover, some of these tools are proprietary and may require payment. This indicates diverse deployment preferences and challenges across CI/CD services, suggesting the need for more integrated and automated solutions to streamline mobile app deployments to popular platforms like the Google Play Store. These inconsistencies in deployment practices can affect reproducibility, maintainability, and contributor onboarding. For example, tool-specific configurations may introduce integration overheads or failures that are hard to debug and replicate. Inconsistent automation strategies can also complicate release planning under the typical rapid development cycles of Android apps. Ultimately, the lack of standardized deployment configurations hinders the uniform adoption of quality assurance practices across projects.
Compare & Contrast with Prior Research Findings (1.4): Gallaba and McIntosh (2018) and Rzig et al. (2022) reported low adoption rates for deployment in Java and Python projects, respectively, with relatively higher rates in Python. Though some CI/CD services like Travis CI offer built-in deployment support (Rzig et al., 2022), the authors found many third-party tools being used for deployment. These tools tend to deploy to cloud services, like AWS or Azure, rather than specific app stores. Moreover, Gallaba and McIntosh (2018) noted that organizations might avoid using the deployment phase in CI due to the potential complexity of its configuration. In the context of Android apps, we also identify several third-party tools used for deployment, but with a noticeable preference for deploying to GitHub, as uploading to app stores is often performed manually. This observation contradicts a prior finding on industrial projects, which advocated for automated deployment over manual methods to accelerate release cycles and minimize errors (Elazhary et al., 2021).
@@
Summary of RQ 1 CI/CD configurations are not standardized and vary in complexity across categories and CI/CD services. Most efforts focus on pipeline setup, comprising over 81% of directives. Testing is performed in nearly half of the projects, with a predominant focus on standard unit testing, whereas deployment is infrequent, occurring in only 9% of the projects. These findings align to some extent with prior research findings on general-purpose projects, though our results contrast in showing lower-than-expected adoption of testing and deployment in Android apps despite their inherently demanding release requirements.
@@ 4.2. RQ2: How do CI/CD configurations evolve in Android apps?
In this RQ, we collect and process the commits in which developers change CI/CD configuration files to understand the complexity of changes, evolution trends, and factors associated with change frequency.
Analyzing complexity and evolution of changes. To analyze the evolution of CI/CD configuration complexity, we segment the data by project and commit period (monthly). To visualize the number of changes (commits) per project across CI/CD services, we use boxplots in which we display commit distribution overall and per CI/CD services, showing variations and outliers. To analyze the evolution of changes across various CI/CD services over the years, we start by identifying the first commit for each project to grasp the initial configuration state. We then calculate the total number of changing commits for each project. These counts are normalized by the total number of active projects adopting CI/CD within a given year, accounting for the newly adopting projects, as follows:
@@
Understanding factors associated with CI/CD change frequency. To understand the factors associated with CI/CD maintenance activity in software projects, by identifying the significant variables associated with the number of configuration commits in the studied projects using linear regression. To do this, we fit an Ordinary Least Squares (OLS) model, a common method for estimating the parameters in a linear regression model, identifying the relationships between variables. To fit our model, we use common project characteristics provided by GitHub about software repositories, as shown in Table 6. Specifically, we use the number of configuration commits as the dependent variable and the other project characteristics as independent variables. We assess the performance of the model using the adjusted $R^{2}$, which measures the proportion of variance in the dependent variable explained by the independent variables. The adjusted $R^{2}$ helps to adjust the obtained $R^{2}$ value for the number of predictors, thus preventing overfitting when using multiple variables. An $R^{2}$ value closer to $1.00$ indicates a better fit. We also evaluate the significance of each factor based on the p-values produced by the model, and identify the directionality of their relationships through coefficient analysis (i.e., either positively associated ($\nearrow$) or negatively associated ($\searrow$)).
@@
Compare & Contrast with Prior Research Findings (2.1): Prior research on general-purpose projects reported lower rates of change in CI configurations for most projects using Travis CI (Hilton et al., 2016a; Zampetti et al., 2021). However, Khatami et al. (2024) found higher change rates of CI/CD pipelines with GitHub Actions. Consistent with the former and in contrast to the latter, our study reveals that Android apps exhibit low change rates across all CI/CD services.
@@
Observation 2.2: The majority of projects (52%) show a decreasing trend in CI/CD changes (commits) over time, with 28% and 20% exhibiting an increasing and steady rate, respectively, of actual configuration lines modified, suggesting simplification or stabilization in CI/CD adoption. Figure 6 shows an overall trend of configuration changes (commits) over time across the different CI/CD services. We observe that GitHub Actions shows a rapid increase in configuration changes, due to shifting towards it after its introduction in 2018. In contrast, Travis CI exhibits a decline after 2017, nearly reaching zero configuration changes after 2020, likely due to the decreased adoption among GitHub projects due to being less popular for open-source projects. CircleCI and GitLab CI/CD demonstrate steady growth peaking in 2018-2019, followed by a steady decrease, indicating moderate user retention, likely amidst the introduction of GitHub Actions. However, the surge in configuration commits for projects adopting GitHub Actions may reflect more than mere popularity; it could also suggest potential instability or complexity, prompting frequent adjustments and updates by developers.
@@
Figure 6. Line plot showing the average number of configuration changes (commits) over time across CI/CD services
With respect to the complexity of configuration changes, we find conflicting trends in the addition and deletion of directives. Specifically, we observe 19% of projects with an increasing number of changes, 42% with a decreasing number, and 41% without significant change. This indicates that many projects are either simplifying their CI/CD setups or maintaining consistent configuration complexity. These trends highlight a dynamic CI/CD landscape, where some projects become more complex or streamlined, while others remain consistent, reflecting differing strategies and stages in project lifecycles.
Moreover, we observe that the frequency of commits has a very weak positive correlation with the number of lines added and deleted, with Pearson coefficients of 0.035 and 0.075, respectively. This indicates that there is no strong link between commit frequency and commit complexity. However, there is a moderate correlation between the trends of committed changes and lines added/deleted over time (Pearson coefficients of 0.27 and 0.28, respectively). This suggests a general increase in development activity and complexity, with commit frequency moderately linked to code changes.
Compare & Contrast with Prior Research Findings (2.2): Gallaba and McIntosh (2018) observed that most Travis CI configuration files rarely change once committed. Though Khatami et al. (2024) observed high volumes of commits modifying GitHub Actions configurations, there was no indication whether such changes increased or decreased over time. Zampetti et al. (2021) support the fact that developers refactor complex CI/CD configurations or use external scripts to simplify their pipelines.
Observation 2.3: CI/CD configurations undergo bi-monthly maintenance (two months, on average), yet many projects have more frequent configuration changes (a median of 2.8 hours), revealing diverse CI/CD maintenance practices. Our analysis of intervals between CI/CD configuration changes across various projects reveals a notable disparity: the average time between changes is approximately 61 days, suggesting a bi-monthly update frequency, while the median of about 2.8 hours indicates that many projects have much more frequent updates. This significant difference, coupled with a high standard deviation of about 84 days, suggests a wide range of maintenance practices in CI/CD configurations. This variation reflects the diverse approaches to CI/CD management among different projects, influenced by factors such as project requirements, team dynamics, and development methodologies.
Compare & Contrast with Prior Research Findings (2.3): While prior studies did not emphasize how frequently developers change CI/CD configurations, they reported that such changes are rare. In addition, prior research observed that some projects experience significantly more modifications to Travis CI configurations than others, aligning with our observations in the context of Android apps across all CI/CD services. Though Ghaleb et al. (2019b, 2022) reported that projects facing build failures or long durations might attempt more frequent configurations to resolve these issues, findings reported by Zampetti et al. (2020) suggest that CI/CD maintenance may be intended to address certain configuration smells or anti-patterns.
@@
Compare & Contrast with Prior Research Findings (2.4): Prior research on Travis CI found that most configuration changes address build job processing nodes or steps (Gallaba and McIntosh, 2018; Zampetti et al., 2021), likely related to setup and build processes. Research on GitHub Actions, however, discovered that most configuration changes deal with testing and other performance issues (Bouzenia and Pradel, 2024). Though Khatami et al. (2024) observed an increase in the adoption of deployment practices in GitHub Actions, we do not observe active maintenance or improvements in the context of Android apps.
Observation 2.5: Active issue tracking, repository size/age, and community engagement are significantly associated with CI/CD configuration maintenance. Table 8 presents the results of the OLS linear regression model, expressed by the modeling coefficients, p-values, significance, and relationship direction. Though our model obtains an adjusted $R^{2}$ value of $0.139$, meaning it explains about 14% of variability in the number of configuration commits, it still offers valuable insights into some of the key factors associated with CI/CD pipeline maintenance activity. We observe that the number of reported issues is significantly associated with CI/CD commit frequency, indicating that developers are likely to update CI/CD pipelines to address issues in their projects. Similarly, repository size and age positively correlate with configuration commits, suggesting larger and older repositories attract more CI/CD activity, possibly due to a larger user base or more bugs/features. Furthermore, we find that repository stars and watchers also positively correlate with commit frequency related to CI/CD configurations, making popular projects more active in addressing issues and deploying app updates, which highlights the community engagement’s role in development activity. Unlike stars and watchers, we find that repository forks are inversely associated with commit frequency. Forks often represent independent development paths, leading to fewer contributions to the original CI/CD pipelines, thus reducing centralized commit activity. Our analysis also explores the relationship between CI/CD services and maintenance activity. We find that integrating GitHub Actions is a significant positive contributor to commit frequency. Other CI/CD services might have positive relationships with commit activity, but lack statistical significance, indicating that evidence is not robust enough to reliably confirm their association. Overall, this analysis highlights the importance of certain factors, such as issue tracking, repository age, community engagement, and adoption of CI/CD practices, in fostering an active development environment, while also identifying potential deterrents to commit activity.
Compare & Contrast with Prior Research Findings (2.5): In regular projects using Travis CI, Felidré et al. (2019) found no correlation between project size and CI change frequency, which contradicts our findings. In terms of project age, Zampetti et al. (2021) found a positive correlation with the frequency of CI changes, aligning with our results. Though prior studies did not explicitly mention associations between CI changes and issue tracking, some studies reported difficulties in troubleshooting build failures (Widder et al., 2019) and adjustments in CI configurations to address these failures (Ghaleb et al., 2022), suggesting a potential link. This suggests that Android apps are distinguished by their unique patterns of CI changes, influenced by factors such as project issue tracking, size, and age.
@@
Summary of RQ 2 Developers regularly change CI/CD pipelines, primarily focusing on setup and build phases, with most projects showing a trend toward simplification over time. While maintenance likely occurs bi-monthly on average, many projects update CI/CD pipelines more often, especially those using GitHub Actions. Increased maintenance activity correlates significantly with active issue tracking, project size/age, and community engagement. Prior research reported varying CI/CD change rates for general-purpose projects depending on the adopted CI/CD service, while change frequency over time was uncertain in those projects.
@@ 4.3. RQ3: What are the common themes in CI/CD maintenance activities in Android apps?
Manual thematic analysis of commit messages. We employ open coding (Corbin and Strauss, 1990) to perform a thorough manual analysis of a statistically significant and stratified random sample of $384$ commit messages, based on a $95\%$ confidence level and a ±$\pm 5$ confidence interval. Our goal in this manual analysis is to identify recurring themes and categories that explain the reasons behind CI/CD configuration changes. Two coders (co-authors of this paper with practical experience in software engineering and CI/CD practices) participate in the manual thematic analysis of the commit messages. They engage in collaborative discussion sessions to ensure consistent interpretation and categorization. Disagreements or ambiguities in coding are resolved through consensus, with the third co-author acting as a mediator to make final decisions. This process ensures the robustness, reliability, and reproducibility of our thematic coding results.
Automated analysis of commit messages. To further validate the results of our manual analysis, we adopt Latent Dirichlet Allocation (LDA) (Blei et al., 2003) as a supplementary step to verify the themes identified manually. Recognizing the importance of parameter tuning in LDA for obtaining meaningful and reliable topic distributions, we follow the recommendations provided by Agrawal et al. (2018). Specifically, we employ LDA with Differential Evolution (LDADE) for parameter optimization, focusing on determining the optimal number of topics and tuning the alpha and beta parameters accordingly. As a result, we obtain alpha and beta values of $0.97$ and $0.73$, respectively, and identify $10$ as the optimal number of topics. Moreover, before performing LDA, we preprocess commit messages to ensure their suitability for topic modeling. First, URLs, email addresses, and default GitHub commit messages are removed. Then, stop words, numbers, and common abbreviations are filtered out, focusing on meaningful content. Lemmatization is also performed for each commit to reduce words to their root forms, and non-English characters are excluded. Also, very short messages (i.e., having fewer than three characters) are discarded as they are often non-informative and do not provide enough context. This thorough preprocessing ensures the text data is clean, standardized, and ready for effective topic modeling with LDA. Finally, we utilize the optimized parameters and processed data to perform LDA topic modeling. For each identified topic, we extract the eight main associated keywords, assign each commit message to the topic with the highest probability of occurrence, calculate the percentage of commit messages that belong to each topic, and assess their relevance to the manually identified themes.
@@
Table 9 lists the common themes of the commit messages identified by our open coding analysis. We describe our identified themes in the following. Our analysis of open codes derived from our manual analysis of commit messages in CI/CD configurations of Android apps reveals several key insights into the current practices and trends in this area.
@@
Compare & Contrast with Prior Research Findings (3.1): A prior study highlighted the challenges developers face in automating CI processes and debugging build failures (Zampetti et al., 2020), while another study showed that configuring CI builds to prevent failures is often related to longer build durations (Ghaleb et al., 2022), reflecting the effort invested in ensuring successful builds. Zampetti et al. (2021) identified several reasons related to changes in Travis-CI configurations. Most importantly, they found that changes to workflow phases and sub-workflows are generally dominant, which aligns with our observation. However, other changes related to documentation and notifications were reported among frequent changes, which is inconsistent with our observation regarding these types of changes. We believe that the themes we identify capture the unique aspects of developing and maintaining Android apps, such as fixing build issues related to dependency on Android components, which differ in release requirements from other applications.
Observation 3.2: Over a third of configuration changes are focused on updating the build environment, tools, and dependencies. Our results show that CI build configurations are more frequently changed for the sake of Environment Update (14.7%), Tool Update (12.6%), and Dependency Update (9.4%). Environment updates are particularly crucial as they demonstrate the necessity of maintaining current environments to align with technological advancements and security protocols. By consistently incorporating these updates, mobile apps remain compatible with the latest technologies, mitigating vulnerabilities inherent in older systems. Moreover, tool and dependency updates highlight the essential need for ongoing management and upgrades of CI/CD pipelines. Regular updates ensure efficient development processes, while managed dependencies keep mobile app components functional and compatible, leading to fewer interruptions and more reliable software performance. While developers should stay updated on CI/CD trends to keep their mobile apps robust, CI/CD service providers should also ensure dependencies are secure (Gruhn et al., 2013) and suggest necessary updates.
@@
Compare & Contrast with Prior Research Findings (3.2): Prior studies (Zampetti et al., 2021; Widder et al., 2019; Khatami et al., 2024; Gallaba and McIntosh, 2018) identified various patterns of changes related to the restructuring of CI configurations. These patterns address multiple aspects like performance, security, and organization, which are often linked, either directly or indirectly, to CI environments and dependencies. Our analysis, however, highlights that changes to the environment and dependencies are particularly prevalent in Android apps, suggesting rapid, ongoing evolutions in the platforms and libraries that support these types of applications.
@@
Observation 3.3: Testing Improvement and Deployment Improvement changes account for 12.8% and 8.4%, respectively. Robust testing and regression in CI/CD pipelines ensure early detection and resolution of bugs, leading to the development of more reliable and high-quality mobile apps. Similarly, deployment ensures rapid and reliable and efficient release cycles, thus reducing the time-to-market for mobile app updates. Despite the importance of frequently maintaining testing and deployment, our results show relatively fewer changes to them compared to other practices, which could be attributed to several factors. These processes might already be well-established and mature within the organization, hence requiring fewer changes and improvements over time, such as enable more features (e.g., add test coverage and add deployment to playstore) or fixing minor relevant issues (e.g. try number two for multiple devices and fix artifact name). In some cases, testing and deployment might not be considered immediate strategic priorities for developers compared to other areas such as feature development, user experience, or performance optimization (e.g., committing to fix tests later). Still, developers are advised to continuously monitor testing and deployment processes, given the critical importance of robust testing and deployment in mobile app deployment, making ongoing improvements and updates are essential to adapt to new technologies and changing requirements. Likewise, researchers should pay more attention to developing techniques that recommend potential co-changes in testing or deployment configuration in response to changes in source code.
Compare & Contrast with Prior Research Findings (3.3): Prior research highlighted the crucial role of testing in CI and identified several configuration patterns related to defining, executing, improving reliability, and optimizing automated test performance in the CI pipeline. However, there was no direct link established between testing and its corresponding changes in CI configurations. Yet, given that developers may use external scripts to run CI testing (Zampetti et al., 2021; Gallaba and McIntosh, 2018), we expect that changes to testing might also be external. Nevertheless, our results still reveal some ratio of changes to testing, though relatively low. For deployment, contrary to our observations, prior research indicated that changes in deployment are somewhat frequent (Gallaba and McIntosh, 2018), which does not seem to apply to Android apps, likely due to the preference for manual deployment.